ADNAN MUHAMMED

Grade 12,

An unknown resistance is to be determined using resistances R1, R and R3. Their correspondins null points are A,B and C. Find which of the above will give the most accurate reading and why?? R=R1 or R2 or R3.


An unknown resistance is to be determined using resistances R1, R and R3. Their correspondins null points are A,B and C. Find which of the above will give the most accurate reading and why??


R=R1 or R2 or R3.

Grade:12

1 Answers

pratham ashish
14 years ago
hi, accod. to wheat stone bridge, null point means that the potential drop across x resistance = potential drop across legth of wire up to the null point. let length r La, Lb Lc resp. for the 3 null points measuring form left hand corner. then, x/La = R1/L-La,(as resistance r parallel) similarly x/Lb =R2/L-LB and so on. so u can know the value of unkwown. accuracy will be in B part as the % error in measuring the length will b minimum.